# CVPT: Cross-Attention help Visual Prompt Tuning adapt visual task

### v1

@@ 3.2 Cross Visual Prompt Tuning
Figure 3: Structure comparison of VPT and CVPT. In which blue represents frozen parameters and orange represents learnable parameters.
Cross Visual Prompt Tuning. We redesign the prompt to better adapt visual tasks and proposed CVPT. Our approach, as illustrated in Fig.3, follows the VPT, the main parameters of the network remain frozen, and only the final classification layer and the prompt are trainable. The key difference is that we allow the prompt token to perform cross-attention with the embedded tokens and the result of cross-attention is added with the embedded tokens as residuals. This operation helps prompts adapt visual tasks a lot, and we demonstrate how significant this improvement is in Sec.4.2. Specifically, for any input $x_{i}$ of a transformer block, the forward flow can be represented as follows:
@@
In CVPT, we only introduce linear computational overhead associated with the number of prompt tokens. It allows CVPT to use a large number of prompt tokens to improve its performance by introducing an acceptable overhead. Furthermore, CVPT preserves the original procedure of self-attention, keeping the complete representation ability of embedded tokens. We demonstrate the improvement over VPT in terms of performance and efficiency in Sec.3.3. Finally, we set embedded tokens as query set and prompt tokens as key-value set, so that we can maintain the unity of the number of channels, allowing the result of cross-attention to be directly summed with the input as a residual term.
@@
Weight-sharing mechanism. The utilization of cross-attention, which requires a large number of learnable parameters (usually $\geq$ 30% model’s parameter number), leads to a major challenge in computational overhead. Therefore, if the parameters of them are tunable, the computational overhead of CVPT will even rival those using full-tuning. Therefore, we introduce the weight-sharing mechanism. Due to the structure of cross-attention equals to that of self-attention, we consider that the weight of self-attention is also instructive for the fine-tuning of cross-attention. Thus, we initialize the weight of cross-attention with the parameters of self-attention when loading checkpoints. It avoids the introduction of a huge number of learnable parameters in cross-attention and keeps the efficiency of our CVPT. We explore the impact of weight-sharing in 4.3 and demonstrate that frozen cross-attention is even more effective than learnable cross-attention.
3.3 Comparison with VPT
Method Number 1 5 10 20 50 100 150 200
VPT 71.0 73.0 73.0 72.8 72.2 69.2 66.0 64.0
CVPT 69.5 73.5 74.0 74.1 74.3 74.5 74.6 74.8
These results show that our CVPT achieves better performance in almost every case except the number of prompts equals 1. As we analyzed in Section 3.1, VPT represents a pool absolute performance on account of the lack of adaptation to visual tasks. Besides, due to the corruption of self-attention between embedded tokens, when given a larger number of prompt tokens, VPT shows significant performance degradation or even crashes. In contrast, our CVPT avoids suffering from these problems. Additionally, its performance improves as the number of prompt tokens increases. All these results above indicate that cross-attention between prompt tokens and embedded tokens helps prompts adapting the visual tasks and instruct the model’s fine-tuning more exactly.
Efficiency improvement. To explore the improvement in efficiency of CVPT, we also recorded the amount of GPU memory occupied by VPT and CVPT during training and testing as well as the total computation of the two when conducting the above experiments, and the results are shown in Fig.4 follows:
Figure 4: The trends of training memory, testing memory, and Flops with the variation in the number of prompt tokens. Where LP represents Linear Probing which only tunes the final classifier linear. We record those data on cifar100 in VTAB-1K, the batch_size is set to 32. Pre-trained model is ViT-B/16.
It can be seen that our CVPT has made significant improvements in efficiency compared to VPT especially given a large amount of prompt tokens. Although it requires slightly more GPU memory during testing compared to full-tuning which is marginal compared to VPT. Additionally, the weight-sharing mechanism allows for targeted optimization in engineering applications, letting cross-attention and self-attention share memory, further widening the efficiency gap with VPT. Moreover, the careful design of CVPT prevents explosive growth in memory and computation as the number of prompts increases. This means we can improve the performance of CVPT by increasing the number of prompts, which is more computationally efficient than other methods.
In summary, our CVPT significantly improves the performance and efficiency of VPT by introducing cross-attention and the weight-sharing mechanism, especially given a larger number of prompts. Therefore, it allows us to introduce more prompts to the prompt-based method in an efficient manner, thus improving its performance. We will demonstrate how much this improvement is and compare it with the SOTA methods in the next section.
4 Experiment
4.1 Experimental settings
  • VTAB-1K. VTAB-1K comprises 19 datasets from different domains, classified into three main categories: the Natural group (natural images captured by standard cameras) cifar svhn caltech pets , the Specialized group (professional images captured by specialized equipment, such as medical and remote sensing images) patchcamelyon erousat , and the Structured group (synthetic images from artificial environments). Each task contains only 1,000 training samples clevr kitti matthey2017dsprites . This is a primary metric for evaluating PEFT’s performance.
  • FGVC. FGVC consists of five fine-grained visual classification benchmarks, including CUB-200-2011 CUB , NABirds birds , Oxford Flowers flower , Stanford-Dogs dogs and Stanford-Cars cars . Unlike VTAB-1K, the datasets in FGVC benchmarks are complete.
  • ADE20K. ADE20K ade20k contains more than 25,000 images and is primarily used for scene perception, parsing, segmentation, multi-object recognition, and semantic understanding. This adaptation is challenging due to the huge gap between the objectives of pretraining and downstream tasks.
Baseline. We primarily use CVPT to compare with the following methods: (1) Full-tuning, (2) Adapter and its improved variants such as LoRA, Adaptformer, RepAdapter, and SPT, and (3) VPT and its variants, including E2VPT, EXPRESS and so on.
Training. We use the ViT-Base-16 model as our main model and AdamW as our optimizer. The other settings and training strategies follow those used in VPT. To avoid extensive hyperparameter search, we only select the number of prompts from [1, 5, 10, 20] for VTAB-1K. Besides, we use single NVIDIA 3090 on VTAB-1K and FGVC benchmark, and use NVIDIA 3090 $\times$ 8 on ADE20k.
4.2 Comparison with the SOTA
VTAB-1K. We compared our method with other baseline methods on the VTAB-1K benchmark. The experimental results are shown in Table.2, where we report the top-1 accuracy of these methods. In the table, we divide the prompt-based methods into one group and the other methods into another group. The bold values in each group represent the best accuracy.
@@ 4.2 Comparison with the SOTA
Linear-probing mae 0 57.6 63.4 85.0 63.2 97.0 86.3 36.6 51.0 78.5 87.5 68.6 74.0 34.3 30.6 33.2 55.4 12.5 20.0 9.6 19.2
Bias bitfit 0.10 65.2 72.8 87.0 59.2 97.5 85.3 59.9 51.4 78.7 91.6 72.9 69.8 61.5 55.6 32.4 55.9 66.6 40.0 15.7 25.1
Adapter adapter 0.15 73.9 69.2 90.1 68.0 98.8 89.9 82.8 54.3 84.0 94.9 81.9 75.5 80.9 65.3 48.6 78.3 74.8 48.5 29.9 41.6
NOAH noah 0.36 75.5 69.6 92.7 70.2 99.1 90.4 86.1 53.7 84.4 95.4 83.9 75.8 82.8 68.9 49.9 81.7 81.8 48.3 32.8 44.2
AdaptFormer adaptformer 0.15 74.7 70.8 91.2 70.5 99.1 90.9 86.6 54.8 83.0 95.8 84.4 76.3 81.9 64.3 49.3 80.3 76.3 45.7 31.7 41.1
LoRA lora 0.29 74.5 67.1 91.4 69.4 98.8 90.4 85.3 54.0 84.9 95.3 84.4 73.6 82.9 69.2 49.8 78.5 75.7 47.1 31.0 44.4
RepAdapter repAdapter 0.23 76.1 72.4 91.6 71.0 99.2 91.4 90.7 55.1 85.3 95.9 84.6 75.9 82.3 68.0 50.4 79.9 80.4 49.2 38.6 41.0
SPT-Adapter spt 0.34 76.2 72.9 93.2 72.5 99.3 91.4 88.8 55.8 86.2 96.1 85.5 75.5 83.0 68.0 51.9 81.2 82.4 51.9 31.7 41.2
SPT-LoRA spt 0.48 76.4 73.5 93.3 72.5 99.3 91.5 87.9 55.5 85.7 96.2 85.9 75.9 84.4 67.6 52.5 82.0 81.0 51.1 30.2 41.3
@@
VPT-Deep vpt 0.53 72.0 78.8 90.8 65.8 98.0 88.3 78.1 49.6 81.8 96.1 83.4 68.4 68.5 60.0 46.5 72.8 73.6 47.9 32.9 37.8
EXPRESS express 0.98 72.9 78.0 89.6 68.8 98.7 88.9 89.1 51.9 84.8 96.2 80.9 74.2 66.5 60.4 46.5 77.6 78.0 49.5 26.1 35.3
DAM-VP dam 2.52 73.1 - - - - - - - - - - - - - - - - - - -
$E^{2}$ e2vpt 0.27 73.9 78.6 89.4 67.8 98.2 88.5 85.3 52.3 87.8 96.1 84.8 73.6 71.7 61.2 47.9 75.8 80.8 48.1 31.7 41.9
CVPT 0.10 76.2 73.0 90.0 73.8 99.2 91.2 90.0 54.4 84.0 96.5 87.2 75.7 78.4 66.7 50.4 81.0 81.5 52.6 33.4 43.3
We first compare our method with other prompt-based methods. The results of our experiments show that our method achieved the best performance among prompt-based methods in 16 out of 19 datasets, significantly outperforming VPT and other VPT-based methods. Notably, CVPT achieves the highest accuracy in all datasets within the structured group, indicating that the addition of cross-attention significantly improves the adaptation of prompts. Therefore, CVPT performs better in those out-of-distribution (OOD) datasets. Additionally, since we use fewer than 20 prompts in VTAB-1K, CVPT requires the lowest number of parameters.
When considering all PEFT methods, we find that on a small dataset like VTAB-1K, almost all mainstream PEFT methods outperformed full-tuning in terms of performance. This suggests that correctly selecting the parameters to fine-tune is crucial. For our CVPT, it shows an impressive performance, which is only 0.2% behind SPT in accuracy while using fewer parameters than SPT, and outperforms the other PEFT methods in performance. This indicates that CVPT reaches SOTA in terms of both performance and parameter count. In particular, compared to other prompt-based methods that show weaknesses, our CVPT deeply explores the potential of prompt-based methods and demonstrates that prompt-based methods can also perform well in the field of visual fine-tuning.
@@
Linear probing mae 85.3 75.9 97.9 86.2 51.3 79.3 0.18
Adapter  adapter 87.1 84.3 98.5 89.8 68.6 85.7 0.41
AdaptFormer  adaptformer 84.7 75.2 97.9 84.7 83.1 85.1 0.37
Bias bitfit 88.4 84.2 98.8 91.2 79.4 88.4 0.28
VPT-Shallow 86.7 78.8 98.4 90.7 68.7 84.6 0.25
VPT-Deep vpt 88.5 84.2 99.0 90.2 83.6 89.1 0.85
DAM-VP dam 87.5 82.1 99.2 92.3 - - -
EXPRESS express 88.3 - 99.0 90.0 80.5 - -
$E^{2}$ e2vpt 88.5 84.2 99.0 90.2 83.6 89.2 0.45
SPT-Adapter spt 89.1 83.3 99.2 91.1 86.2 89.8 0.41
SPT-LoRA spt 88.6 83.4 99.5 91.4 87.3 90.1 0.48
CVPT 89.7 86.1 99.3 91.4 84.9 90.3 0.79
Similar to the results on VTAB-1K, our approach substantially outperforms other prompt-based methods on FGVC benchmark. Additionally, it surpasses SPT and other adapter-based methods to achieve the best performance. This suggests that CVPT exhibits better performance on relatively large datasets like FGVC, which proves the adaptability of CVPT to the increasing scale of data in the future.
ADE20K. Finally, we apply CVPT to SETRsetr on the ADE20K dataset to explore its performance on downstream tasks of semantic segmentation. The results are shown in Table.4 below:
Table 4: Results of ADE20K datasets with ViT-L models. We report "mIoU-SS" and "mIoU-Ms" which denote single-scale and multi-scale, respectively
@@
Bias bitfit 13.46 43.40 45.33
VPT vpt 13.43 42.11 44.06
RepAdapter repAdapter 13.82 44.44 46.71
SPT-Adapter spt 14.60 45.20 47.20
SPT-LoRA spt 14.60 45.40 47.50
@@
This task is quite challenging because of the huge distribution gap between pre-training datasets and downstream tasks. In this situation, our CVPT shows a 1.7% enhancement of "mIoU-SS" over the VPT with the same number of prompts. If we use 200 prompts for fine-tuning, CVPT represents a significant improvement over the other PEFT methods. This fully demonstrates the adaptation of CVPT to OOD datasets. Besides, due to our optimization of the deployment, even though the number of learnable parameters increases by 4 million, our memory usage and training time increase by less than 10% compared to linear probing and less than 5% compared to it when using 10 prompts during training.

### v2

@@ 4.1 Cross-Attention
4.2 Cross Visual Prompt Tuning
We propose CVPT, which redesigns the deployment of prompts to better adapt visual tasks. Our approach, as illustrated in Fig.3, follows the VPT, the main parameters of the network remain frozen, and only the final classification layer and the prompt are trainable. The key difference is that we allow the prompt token to perform cross-attention with the embedded tokens and the result of cross-attention is added with the embedded tokens as residuals. This operation allows us to decouple the self-attention between prompts and embedded tokens, and more efficiently reestablish the connection between the prompts and embedded tokens. We demonstrate how significant this improvement is in Sec.5.2. Specifically, for any input $x_{i}$ of a transformer block, the forward flow can be represented as follows:
@@ 4.2 Cross Visual Prompt Tuning
In CVPT, we only introduce linear computational overhead associated with the number of prompt tokens. It allows CVPT to use a large number of prompt tokens to improve its performance by introducing an acceptable overhead. Furthermore, CVPT preserves the original procedure of self-attention, keeping the complete representation ability of embedded tokens. We demonstrate the improvement over VPT in terms of performance and efficiency in Sec.5.3. Finally, we set embedded tokens as query sets and prompt tokens as key-value sets, so that the consistency of the dimension can be maintained, allowing the result of cross-attention to be directly summed with the input as a residual term.
4.3 Weight-sharing mechanism
@@
The utilization of cross-attention, which requires a large number of learnable parameters, leads to a major challenge in computational overhead. Therefore, if the parameters of them are tunable, the computational overhead of CVPT will even rival those using full-tuning. Therefore, we introduce the weight-sharing mechanism. Due to the structure of cross-attention equal to that of self-attention, we consider that the weight of self-attention is also instructive for the fine-tuning of cross-attention. Thus, we initialize the weight of cross-attention with the parameters of self-attention when loading checkpoints. It avoids the introduction of a huge number of learnable parameters in cross-attention and keeps the efficiency of our CVPT. We explore the impact of weight-sharing in Sec.5.4 and demonstrate its effectiveness and efficiency.
@@
Linear-probing [21] 0 57.6 63.4 85.0 63.2 97.0 86.3 36.6 51.0 78.5 87.5 68.6 74.0 34.3 30.6 33.2 55.4 12.5 20.0 9.6 19.2
Bias [55] 0.10 65.2 72.8 87.0 59.2 97.5 85.3 59.9 51.4 78.7 91.6 72.9 69.8 61.5 55.6 32.4 55.9 66.6 40.0 15.7 25.1
Adapter [24] 0.15 73.9 69.2 90.1 68.0 98.8 89.9 82.8 54.3 84.0 94.9 81.9 75.5 80.9 65.3 48.6 78.3 74.8 48.5 29.9 41.6
AdaptFormer [5] 0.15 74.7 70.8 91.2 70.5 99.1 90.9 86.6 54.8 83.0 95.8 84.4 76.3 81.9 64.3 49.3 80.3 76.3 45.7 31.7 41.1
LoRA [25] 0.29 74.5 67.1 91.4 69.4 98.8 90.4 85.3 54.0 84.9 95.3 84.4 73.6 82.9 69.2 49.8 78.5 75.7 47.1 31.0 44.4
RepAdapter [36] 0.23 76.1 72.4 91.6 71.0 99.2 91.4 90.7 55.1 85.3 95.9 84.6 75.9 82.3 68.0 50.4 79.9 80.4 49.2 38.6 41.0
SPT-LoRA [19] 0.48 76.4 73.5 93.3 72.5 99.3 91.5 87.9 55.5 85.7 96.2 85.9 75.9 84.4 67.6 52.5 82.0 81.0 51.1 30.2 41.3
RLRR [12] 0.33 76.8 75.6 92.4 72.9 99.3 91.5 89.8 57.0 86.8 95.2 85.3 75.9 79.7 64.2 53.9 82.1 83.9 53.7 33.4 43.6
DM-LoRA [13] 0.29 77.0 74.0 90.7 73.9 99.3 92.2 91.1 56.4 85.6 96.5 87.0 76.1 83.5 69.9 52.0 81.6 80.2 50.2 36.1 43.1
@@
VPT-Deep [27] 0.53 72.0 78.8 90.8 65.8 98.0 88.3 78.1 49.6 81.8 96.1 83.4 68.4 68.5 60.0 46.5 72.8 73.6 47.9 32.9 37.8
EXPRESS [9] 0.98 72.9 78.0 89.6 68.8 98.7 88.9 89.1 51.9 84.8 96.2 80.9 74.2 66.5 60.4 46.5 77.6 78.0 49.5 26.1 35.3
DAM-VP [26] 2.52 73.1 - - - - - - - - - - - - - - - - - - -
$E^{2}$ [18] 0.27 73.9 78.6 89.4 67.8 98.2 88.5 85.3 52.3 87.8 96.1 84.8 73.6 71.7 61.2 47.9 75.8 80.8 48.1 31.7 41.9
VAPT [32] 0.25 75.3 80.8 91.9 69.7 98.8 89.2 86.7 52.9 84.4 96.5 85.1 74.5 74.8 63.6 50.0 77.2 86.1 48.3 33.8 40.9
VFPT [56] 0.48 75.5 80.7 91.4 69.4 99.3 90.3 85.6 52.7 83.5 96.5 84.4 75.4 75.8 63.2 48.3 79.3 81.5 56.0 34.1 43.4
[…]
CVPT 0.39 77.2 81.5 91.5 74.0 99.2 91.4 90.7 54.5 85.8 96.5 87.6 75.8 79.0 67.2 50.6 82.7 81.5 53.0 34.4 45.3
5 Experiment
5.1 Experimental settings
  • VTAB-1K. VTAB-1K comprises 19 datasets from different domains, classified into three main categories: the Natural group (natural images captured by standard cameras) [30][38][14][40], the Specialized group (professional images captured by specialized equipment, such as medical and remote sensing images) [48][23], and the Structured group (synthetic images from artificial environments). Each task contains only 1,000 training samples [28][16][37]. This is a primary metric for evaluating PEFT’s performance.
  • FGVC. FGVC consists of five fine-grained visual classification benchmarks, including CUB-200-2011 [49], NABirds [46], Oxford Flowers [39], Stanford-Dogs [29] and Stanford-Cars [15]. Unlike VTAB-1K, the datasets in FGVC benchmarks are complete.
  • ADE20K. ADE20K [59] contains more than 25,000 images and is primarily used for scene perception, parsing, segmentation, multi-object recognition, and semantic understanding. This adaptation is challenging due to the huge gap between the objectives of pretraining and downstream tasks.
Baseline. We primarily use CVPT to compare with the following methods: (1) Full-tuning and Linear Probing, (2) Adapter and its variants, and (3) VPT and its variants.
Training. We use the ViT-Base-16 model as our main model and AdamW as our optimizer. The other settings and training strategies follow those used in VPT. We set the number of prompts from [1, 5, 10, 20, 50, 100, 200] for VTAB-1K (consistent with VPT). Besides, we use a single NVIDIA 3090 on VTAB-1K and FGVC benchmark and use NVIDIA 3090 $\times$ 8 on ADE20k.
5.2 Comparison with other PEFT methods
VTAB-1K. We compared CVPT with other baselines on the VTAB-1K benchmark. The experimental results are shown in Table.1, where we report the top-1 accuracy of these methods. In the table, we divide the prompt-based methods into one group and the other methods into another group. The bold values in each group represent the best accuracy.
[…]
When considering all PEFT methods, we find that on a small dataset like VTAB-1K, almost all mainstream PEFT methods outperformed full-tuning in terms of performance. This suggests that correctly selecting the parameters to fine-tune is crucial. Our CVPT, shows an impressive performance, over 0.2 in accuracy than DM-LoRA [13] and 0.4 than RLRR [12], outperforming the other PEFT methods in performance in VTAB-1K. Notably, given the previously observed limitations of prompt-based methods in terms of performance and efficiency, our CVPT deeply explores the potential of prompt-based methods and demonstrates that prompt-based approaches can be competitive with state-of-the-art adapter-based methods.
@@ 5.2 Comparison with other PEFT methods
Linear probing [21] 85.3 75.9 97.9 86.2 51.3 79.3 0.18
Adapter  [24] 87.1 84.3 98.5 89.8 68.6 85.7 0.41
AdaptFormer  [5] 84.7 75.2 97.9 84.7 83.1 85.1 0.37
LoRA [25] 88.3 85.6 99.2 91.0 83.2 89.5 0.44
SPT-LoRA [19] 88.6 83.4 99.5 91.4 87.3 90.1 0.48
RLRR [12] 89.3 84.7 99.5 92.0 87.0 90.4 0.47
DMLoRA [13] 89.8 86.6 99.5 91.8 85.7 90.7 0.47
VPT-Shallow [27] 86.7 78.8 98.4 90.7 68.7 84.6 0.25
VPT-Deep [27] 88.5 84.2 99.0 90.2 83.6 89.1 0.85
DAM-VP [26] 87.5 82.1 99.2 92.3 - - -
EXPRESS [9] 88.3 - 99.0 90.0 80.5 - -
$E^{2}$ [18] 88.5 84.2 99.0 90.2 83.6 89.2 0.45
VAPT [32] 89.7 84.6 99.1 91.7 82.8 89.6 0.67
VFPT [56] 88.7 84.5 99.1 90.4 83.6 89.2 0.85
[…]
CVPT 89.9 86.5 99.3 91.7 85.2 90.5 0.77
Similar to the results on VTAB-1K, our approach substantially outperforms other prompt-based methods and achieves the best results in 4 out of 5 datasets in FGVC. It is only behind 0.2% than DMLoRA [13] and outperforms other methods. This demonstrates CVPT’s generalization and adaptability to the increasing scale of data in the future.
ADE20K. Finally, we apply CVPT to SETR[58] on the ADE20K dataset to explore its performance on semantic segmentation tasks. The results are shown in Table.3.
Table 3: Results of ADE20K datasets with ViT-L. "mIoU-SS" and "mIoU-Ms" denote single-scale and multi-scale, respectively.
@@
Bias [55] 13.46 43.40 45.33
VPT [27] 13.43 42.11 44.06
VPT+Bias [27] 15.79 44.04 45.63
RepAdapter [36] 13.82 44.44 46.71
SPT-LoRA [19] 14.60 45.40 47.50
@@
This task is quite challenging because of the huge distribution gap between pre-training datasets and downstream tasks. In this situation, our CVPT shows a 1.7% enhancement of "mIoU-SS" over the VPT with the same number of prompts. If we use 200 prompts for fine-tuning, CVPT represents a significant improvement over the other PEFT methods. This fully demonstrates the adaptation of CVPT to OOD datasets. Besides, due to our optimization of the deployment, even though the number of learnable parameters increases by 4 million, our memory usage and training time increase by less than 20% compared to linear probing and less than 10% compared to it when using 10 prompts during training.
5.3 Comparison with VPT
Method Number 1 5 10 20 50 100 150 200
VPT 71.0 73.0 73.0 72.8 72.2 69.2 66.0 64.0
CVPT 69.5 73.5 74.0 74.1 74.3 74.5 74.6 74.8
These results show that our CVPT achieves better performance in almost every case except the number of prompts equals 1. As we analyzed in Sec.3, due to the destruction of self-attention between embedded tokens, when given a larger number of prompt tokens, VPT shows significant performance degradation or even crashes. In contrast, our CVPT avoids suffering from these problems. Additionally, its performance improves as the number of prompt tokens increases. We selected CIFAR[30], DTD, and SUN397 to investigate the reasons for the performance degradation in VPT, and the results are shown in Table.5.
Table 5: Performance comparisons With VPT and CVPT on CIFAR, DTD, and SUN397, whose feature distribution is similar to pretraining dataset.
Dataset Method Number 1 10 20 50 100 200
CIFAR VPT 65.2 64.9 63.6 60.3 57.5 35.7
CVPT 70.2 72.4 72.0 72.6 71.9 72.1
DTD VPT 68.8 66.1 65.9 63.4 61.9 59.5
CVPT 70.7 72.4 72.0 72.6 71.9 72.1
SUN397 VPT 52.6 47.6 46.8 43.6 34.4 27.5
CVPT 52.3 54.4 54.2 54.1 53.9 54.0
Han et al. [17] think pretrained parameters play a pivotal role in capturing general features, while the added learnable parameters are important for potentially encoding task information in the context of transfer learning. This indicates that the destruction of self-attention between embedded tokens weakens the ability to recognize general features, and significantly leads to performance collapse on datasets with feature distributions similar to those of the pretraining data. Meanwhile, it can be observed that the performance of our CVPT does not fluctuate significantly with variations in the number of prompts. This demonstrates the advantage of preserving the complete self-attention.
Figure 4: The trends of training memory, testing memory, and Flops with the variation in the number of prompt tokens. Where LP represents Linear Probing which only tunes the final classifier linear. The batch_size is set to 32. Pre-trained model is ViT-B/16.
Efficiency improvement. To explore the improvement in the efficiency of CVPT, we also recorded the amount of GPU memory occupied by VPT and CVPT during training and testing as well as the total computation of the two when conducting the above experiments, and the results are shown in Fig.4 follows.
The results reveal that our CVPT has made significant efficiency improvements compared to VPT especially given a large amount of prompt tokens. Additionally, the weight-sharing mechanism allows for targeted optimization in engineering applications, letting cross-attention and self-attention share memory, further widening the efficiency gap with VPT. Moreover, the careful design of CVPT prevents explosive growth in memory and computation as the number of prompts increases. This means we can improve the performance of CVPT by increasing the number of prompts more computationally and efficiently than before.
In summary, our CVPT significantly improves the performance and efficiency of VPT by introducing cross-attention and the weight-sharing mechanism, especially given a larger number of prompts. Therefore, it allows us to introduce more prompts to the prompt-based method in an efficient manner, thus improving its performance in OOD datasets.